# Global Prediction of COVID-19 Variant Emergence Using Dynamics-Informed Graph Neural Networks

### v2

@@ III-C Dynamics-based Model
Suppose we select $t_{0}$ as the time at which $p^{X}_{i}(t_{0})$ is large enough such that the community transmission dictates the dynamics in the region $Y$ rather than importations and so, both Equations 8 & 9 hold for $t\geq t_{0}$. Suppose that the threshold for prevalence in region $Y$ for community transmission to dominate is $\theta_{Y}$ and that for region $X$ is $\theta_{X}$. We assume that $\theta_{X}$ and $\theta_{Y}$ are independent of the variant. Then, by setting $t=t_{0}$ in equations 8 & 9, we get:
@@
Therefore, both $C_{Y}^{\prime}$ $C_{X}^{\prime}$ are variant independent. Let $t^{\theta}_{Y}$ and $t^{\theta}_{X}$ be the times at which variant $i$ reach target prevalence of $\theta$ in regions $Y$ and $X$, respectively. We want to find $\tau=t^{\theta}_{Y}-t^{\theta}_{X}$. From equations 8 and 9:
@@
We then use the calculated weights from our linear models in equation 14 and plug them into equation 13 to find the $\tau$ value between two regions. This is a pairwise model that estimates the delay between two regions rather than providing a delay from the current date. Algorithm 1 shows how we compute the date at which a variant arrives in a given region $Y$, we calculate the median of the outputs from all pairwise models from region $X_{p}$ to region $Y$. We evaluate the performance of all the models $(X_{p},Y)$ until the current date, and pick the top three source regions among $\{X_{1},X_{2},\dots\}$. The median of these selected models is then taken to represent the delay at time $t$ for the new variant.
@@
1:for all $X_{p}$ in $\{X_{1},X_{2},\dots\}_{t}$ do
2:     Calculate and store delay predictions for $(X_{p},Y)$ until the current date, $t$.
@@
4:     Select top 3 $X_{p}$ with the best performance at $t-1$
@@ III-D1 Graph Construction
The graph construction is illustrated in Figure 3. First, we represent each country as a node in the graph and edges in the graph are established to represent relationships between the countries. There are a total of 87 countries in which at least one variant appears and becomes dominant in the CoVariants dataset [covariants]. While not shown in Figure 3, self-loops are included to account for internal transmissions. Next, we consider temporal aspect of the problem. The graph may evolve over time to capture changing relationships or influences among countries. This is introduced into our graph as dynamic edge weights based on border control data in the OxCGRT dataset [hale2021global]. Section IV-A provides more details on this dataset and its pre-processing.
@@
To obtain the full graph, we create multiple copies of the above graph representing regions and inter-connectivity – one per variant. This approach allows us to leverage variant-specific dynamics as our node features. For each node $c$, the considered features consist a time series of the log of prevalence ratios $r$ of a present variant $i$ as defined in equation 2, for $T$ time steps. This results in a vector $r_{c,i}$ and the corresponding $S_{i}$ value of this variant. These components are then concatenated to form each node’s final feature vector $[r_{c,i}|S_{i}]$. The selection of these features was informed by our derivation in Equation 5, which shows that, under some assumptions, there exists a simple linear relation between the two. We hypothesize that these features simplify the underlying patterns that are then more easily learned by a GNN that may complement the violation of the assumption of the dynamics-based model.
III-D2 Temporal Encoded Graph Convolutional Network (T-GCN)
@@
In order to extract both temporal and geographical information of the data, we propose a hybrid architecture summarized in Figure 4. In this model, we employ a simple Gated Recurrent Unit (GRU) [cho2014learning] to extract temporal information of the variant transmissions in $T$ time steps. We temporarily remove the growth rate $S_{i}$, which is not time-dependent, from node features and embed the 1-dimensional (1D) time-dependent features (prevalence ratios) into a 2D latent feature vector. Note that in this configuration, we embed the features separately for each country, treating the nodes as distinct input samples grouped into a single batch, resulting in a 3D latent feature vector. The 3D latent feature vector is then flattened to obtain a final 2D embedding, which serves as input for a one-layer Graph Convolutional Network (GCN). The Leaky ReLU activation function introduces non-linearity after the GCN layer. Additionally, two GraphNorm[GraphNorm] layers are applied after the GRU and GCN layers for normalization, and dropout is incorporated after the GCN layer for regularization. Finally, the embedded features are fed into a Fully Connected (FC) layer to predict delay and emergence for each country.
III-D3 Dynamics-Informed Loss GCN (DIL-GCN)
In this approach, we explicitly guide our GCN, described in Section IV-D, to make predictions based on dynamics, specifically for the delay prediction task. This guidance is achieved by incorporating information from our dynamics-based model. Given that our GCN is trained using a Mean Squared Error (MSE), the loss function is:
MSE $=\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}$ (15)
where N is the number of countries, $\hat{y}_{i}$ is our models output, and $y_{i}$ is the ground truth. We modify this loss function to steer our GCN towards dynamic predictions inspired by the dynamics-based model. This technique is typically used in PINNs [raissi2019physics]. The adjusted loss function is given by:
@@ III-D3 Dynamics-Informed Loss GCN (DIL-GCN)
where $p$ is a hyperparameter ranging from 0 to 1, signifying the influence of $\dot{y}_{i}$, the output from the dynamics-based model. This modification ensures that, during training, the model is penalized more when its prediction deviates from the linear model’s prediction, providing a form of dynamics-informed regularization. The best $p$ value was found to be 0.1 by validating on both logarithmic and linear ranges of $p$.
Figure 4: T-GCN Architecture. The prevalence ratio for each time step is encoded into a 2D latent feature vector by the GRU. Then, the feature vectors are flattened and the growth rate is concatenated to them. Our batch size of $N$ = # of countries.
@@ III-E Training Procedure
3:     model_initial_weights $\leftarrow$ random
@@
9:         model $\leftarrow$ initialize with $model\_initial\_weights$
10:         epochs $\leftarrow$ 100
11:         early_St $\leftarrow$ Early Stopper(patience=3)
@@
13:         best_model_weights $\leftarrow early\_St$
@@
All training and validation of the models were performed retrospectively, as outlined in Algorithm 2. This means that the model is trained and validated at each time step using only the ”observed” data available up to that time step, denoted as $d$. For each variant, the algorithm iterates through all the biweekly data associated with that variant, training the model only on the weeks that have already passed (i.e., the observed data). It is important to note that the pre-processing is also done retrospectively, accounting for any smoothing, calculation of the $S$ values, or interpolation in the case of the dynamics-based model. This ensures that the approach is applicable in a prospective setting, where no data from the future is available.
The train/validation split is done temporally to ensure that validation is performed only on the most recently observed data. This approach aims to achieve a better fit to the latest data, enhancing the model’s predictive performance. An Early Stopper monitors the validation loss and halts the training when overfitting is detected, signified by a constant and significant increase in the validation loss. The Early Stopper saves the best-performing weights, which are then reused in the next iteration as a form of transfer learning. This practice eliminates the need for the model to start training from scratch, leveraging knowledge gained from previous iterations.
While Algorithm 2 shows the general approach taken for the training, problem-specific aspects are discussed below.
III-E1 Delay Prediction
Since this is a regression task, we use MSE as the loss function (Equation 15). Another crucial aspect is how to handle data related to variants that either do not appear in a given country or appear but fail to become dominant, i.e. rapidly diminish. In such cases, where regression targets would be undefined (potentially infinite), we address this challenge by creating a mask. This mask is employed to conceal nodes corresponding to these specific variants/country pairs, ensuring they do not contribute to our training loss. Essentially, our graph models are not trained on these nodes.
III-E2 Emergence Prediction
For the problem of emergence prediction, the chosen loss function is the Binary Cross Entropy (BCE) Loss. In addition to this, we note that when a new variant appears, it only reaches a maximum of 38 countries out of the 87, resulting in a large class imbalance. To account for the class imbalances, class weights are applied to the loss function, providing a balanced learning approach. These weights were calculated using the Inverse Class Frequency method [classBalance].
@@ IV-A1 Covariant Data
CoVariants [covariants] is our primary dataset which provides the current state of variants and mutations of interest for SARS-CoV-2. This data is enabled by GISAID [gisaid] and ranges from August 2020 through October 2023, with a bi-weekly resolution for 36 variants starting from 20A.EU1 to 23F.Omicron, which is current as of this work. CoVariants follows the NextStrain Clade schema [nextstrain-clade], in which variants can descend from other variants.
@@ IV-A2 Border Restrictions Data
We considered the border restrictions data provided in the OxCGRT [hale2021global] dataset, with details on what level of border controls were enacted by governments in real-time from Jan 1, 2020, to Jan 8, 2023. OxCGRT measured the variations in government responses using their COVID-19 Government Response Stringency Index, which is a simple additive score of nine indicators ranging from school closures to vaccination policies. Of these indicators, we picked international travel closure controls since they best represent the goal of this paper in capturing the emergence and cross-border transmission of variants. These values are encoded in an ordinal scale as follows: 0 - no measures, 1 - screening, 2 - quarantine on high risk region, 3 - ban on high risk regions, 4 - total border closure. The recorded values were then scaled down to a range between 0.1 and 1, with 1 indicating no border restrictions and 0.1 representing complete border closure. Note that 0.1 was chosen over 0, acknowledging the possibility of some mobility between connecting countries due to necessary trade or border crossings via open land routes. Furthermore, given that our CoVariants data extends beyond January 2023, we extend OxCGRT by making the assumption that all border restrictions have remained unchanged since Jan 8, 2023.
@@ IV-A3 Country Graphs
We considered two graphs representing the interconnection between countries. This dataset depicts country neighborhoods according to the Correlates of War (CoW) database. Countries were also deemed neighbors if they share a land or river border or are located within 24 miles of each other across bodies of water. Adjacency: We used the country adjacency dataset [countryAdjacencyData], which offers a list of countries and their neighboring countries. OpenFlights: This is a collection of regular flights along with their routes mapped from airports around the world [OpenFlights2023]. Given the scope of the current work, we primarily focus on airports and the flight route data, and was mainly used during ablation studies. Airports data is a UTF-8 encoded collection of over 14,000 airports, train stations and ferry terminals across the world as of January 2017. It consists of information like city/country, IATA codes, geographic coordinates and the type of the airport (air terminal, ferry station, etc.). Routes data provides information about 67663 directional routes between 3321 airports on 548 airlines. This dataset was used in an ablation study to validate the effectiveness of utilizing these flight routes as edges in our graph creation process.
@@ IV-B Benchmark Pipeline
As previously mentioned, we are the first to attempt this type of problem as far as we know. In Figure 5, we present the comprehensive pipeline employed for the retrospective training, validation, and testing of diverse models. The design of this pipeline prioritizes user-friendliness, allowing seamless integration of new PyTorch models or direct utilization of scikit-learn models through a configuration file. Users only need to specify whether the model requires the use of graph data and the nature of the task, whether it involves Delay Prediction or Emergence Prediction, leaving the pipeline to handle the rest. Additionally, the data pre-processing is encapsulated in a separate function, providing users with the flexibility to make any necessary modifications. This approach facilitated the training of various benchmark models on our dataset and extends an accessible framework for users to experiment with their own models.
@@
In the context of delay prediction, considering the biweekly data, the target is consistently a multiple of 14 days. Therefore, the output of our models is always rounded up to the nearest multiple of 14. Furthermore, evaluations exclude countries where a variant has already become dominant or never achieves dominance. To measure performance, we define Mean/Median Absolute Errors (MAE/MedAE) as follows:
@@
In the context of emergence prediction, our objective is to retrospectively predict whether a given variant will become dominant in each country. Given the presence of class imbalance in our data, we employ the Macro F1-score as an evaluation metric, which considers both precision and recall across all classes, providing a balanced assessment [opitz2019macro, lipton2014optimal].
IV-C Pre-processing
The initial pre-processing stage focused on aligning the country names across diverse datasets, with an emphasis on inclusion based on data availability. Only countries present in all datasets were retained, ensuring a standardized set for subsequent analyses. A specific adjustment was necessary for the OpenFlights dataset, involving the mapping of city routes to their respective country routes. Leveraging the provided airport data, which contains comprehensive city and country details, facilitated this mapping process. The transformation ensured that subsequent analyses operated consistently at the country level.
For each country, the prevalence of each variant, defined in equation 1, was calculated at each timestep. Variants with a prevalence of less than 5% were disregarded in our analysis. To calculate the prevalence ratios, the two variants with the highest prevalence were identified for each timestep. For each variant on that particular day, the prevalence ratio was calculated using equation 2. Finally, variants that did not reach a prevalence ratio of 0.2 or those that were present in less than three countries were filtered out. These steps ensured that only variants with a significant presence and wide geographic distribution were included in the subsequent analysis. All subsequent pre-processing steps were conducted retrospectively. This implies that as time progresses, more data is used for computing the global growth rates ($S$) and more data is available for training. Notably, for our dynamics-based model, this retrospective process also involved interpolating biweekly data into a daily resolution using a cubic smoothing spline. This adjustment was deemed necessary due to the transient nature of variants. The biweekly data has a resolution that is too coarse to adequately capture the linear relationships between the delay in variant transmission from one country to another and its prevalence in the latter.
IV-D Benchmark Models

### v3

@@ III-C Dynamics-based Model
Suppose we select $t_{0}$ as the time at which $p^{X}_{i}(t_{0})$ is large enough such that the community transmission dictates the dynamics in the region $Y$ rather than importations and so, both Equations 6 & 7 hold for $t\geq t_{0}$. Suppose that the threshold for prevalence in region $Y$ for community transmission to dominate is $\theta_{Y}$ and that for region $X$ is $\theta_{X}$. We assume that $\theta_{X}$ and $\theta_{Y}$ are independent of the variant. Then, by setting $t=t_{0}$ in equations 6 & 7, we get:
@@
Therefore, both $C_{Y}^{\prime}$ $C_{X}^{\prime}$ are variant independent. Let $t^{\theta}_{Y}$ and $t^{\theta}_{X}$ be the times at which variant $i$ reach target prevalence of $\theta$ in regions $Y$ and $X$, respectively. We want to find $\tau=t^{\theta}_{Y}-t^{\theta}_{X}$. From equations 6 and 7:
@@
We then use the calculated weights from our linear models in equation 12 and plug them into equation 11 to find the $\tau$ value between two regions. This is a pairwise linear model that estimates the delay between two regions rather than providing a delay from the current date. Algorithm 1 details how we compute the date at which a variant arrives in a region $Y$. We calculate the median of the outputs from all pairwise models from regions $X_{q}$ to region $Y$. We evaluate the performance of all the models $(X_{q},Y)$ until the current date and select the top three performing source regions among ${X_{1},X_{2},\dots}$. Notably, best performance here refers to the lowest error in predicting the delay. We identify three source regions that produced the lowest error in predicting the delays at $t-1$. Then, we take the median of the predictions produced by these selected models at time $t$ which represents the delay at time $t$ for the new variant to appear in region $Y$.
@@
1:  for $X_{q}$ in $\{X_{1},X_{2},\dots\}_{t}$ do
2:     Calculate and store delay predictions for $(X_{q},Y)$ until the current date, $t$.
@@
4:     Select top 3 $X_{q}$ with the best performance at $t-1$
@@ III-D Dynamics-Informed GNN
Key Idea: We observe that global growth rates $S_{i}$ play a crucial role in the dynamics. Furthermore, based on Equation 3, under some assumptions, the logarithm of the prevalence ratio $r_{ij}$ grows linearly with time. We hypothesize that using $\ln(r_{ij})$ and $S$ as features simplify the underlying patterns to be learned by an ML algorithm. An ML algorithm may also complement any violations of the assumptions made in the dynamics-based model.
@@ III-D1 Graph Construction
The graph construction is illustrated in Figure 3. First, we represent each country as a node in the graph, and edges are established to represent relationships between the countries. There are 87 countries in which at least one variant appears and becomes dominant in the CoVariants dataset [11]. The dataset provides the state of variants and mutations of interest for COVID-19 from August 2020 through October 2023. It has a bi-weekly resolution for 36 variants starting from 20A.EU1 to 23F.Omicron. While not shown in Figure 3, self-loops are included to account for internal transmissions.
Next, we consider the temporal aspect of the problem. The graph may evolve over time to capture changing relationships or influences among countries. This is introduced into our graph as dynamic edge weights based on border control data in the “OxCGRT” dataset [31]. This dataset is from January 2020 to January 2023 measuring the variations in government responses using their COVID-19 Government Response Stringency Index, which is a simple additive score of nine indicators ranging from school closures to vaccination policies. Of these indicators, we picked international travel closure controls since they best capture the emergence and cross-border transmission of variants. These values are encoded in an ordinal scale as follows: 0 - no measures, 1 - screening, 2 - quarantine on high-risk regions, 3 - ban on high-risk regions, 4 - total border closure. The recorded values were then scaled down to a range between 0.1 and 1, with 1 indicating no border restrictions and 0.1 representing complete border closure. Note that 0.1 was chosen over 0, acknowledging the possibility of some mobility between connecting countries due to necessary trade or border crossings via open land routes. Given that our CoVariants data extends beyond January 2023, we extend OxCGRT by assuming that all border restrictions have remained unchanged since Jan 8, 2023.
@@
To obtain the full graph, we create multiple copies of the above graph representing regions and inter-connectivity – one per variant at a given timestamp $t$. This approach allows us to leverage variant-specific dynamics as our node features. At $t$, each node $c$ has features consisting of a time series of the logarithm of prevalence ratios $r_{ij}$ of a present variant $i$ and the most prevalent variant $j\neq i$, spanning $T$ time steps. This results in a vector $\vec{r}_{c,i}=\left[\ln\left(r_{ij}^{t-T}\right),\ln\left(r_{ij}^{t-T+1}\right),\dots,\ln\left(r_{ij}^{t}\right)\right]$. The corresponding $S_{i}$ value of the variant $i$ is then concatenated to form each node’s final feature vector $[\vec{r}_{c,i}|S_{i}]$.
III-D2 Feature Augmented Graph Convolutional Network (FA-GCN)
Figure 4: FA-GCN Architecture. The prevalence ratio for each time step is encoded into a 2D latent feature vector by the GRU. Then, the feature vectors are flattened and the growth rate is concatenated to them. Our batch size of $N$ = # of countries.
@@ III-D2 Feature Augmented Graph Convolutional Network (FA-GCN)
To extract both temporal and geographical information of the data, we propose a hybrid architecture summarized in Figure 4. In this model, we employ a simple Gated Recurrent Unit (GRU) [32] to extract temporal information of the variant transmissions in $T$ time steps. We temporarily remove the growth rate $S_{i}$ (which is not time-dependent) from node features and embed the 1-dimensional (1D) time-dependent features (prevalence ratios) into a 2D latent feature vector. Note that in this configuration, we embed the features separately for each country, treating the nodes as distinct input samples grouped into a single batch, resulting in a 3D latent feature vector. The 3D latent feature vector is then flattened to obtain a final 2D embedding, which serves as input for a one-layer Graph Convolutional Network (GCN). The Leaky ReLU activation function introduces non-linearity after the GCN layer. Additionally, two GraphNorm[33] layers are applied after the GRU and GCN layers for normalization, and dropout is incorporated after the GCN layer for regularization. Finally, the embedded features are fed into a Fully Connected (FC) layer to predict delay until emergence for each country.
III-D3 Physics Informed GCN (PI-GCN)
For comparison, we use a PINN-inspired [10] approach to make predictions based on dynamics for the Delay Prediction task. This model utilizes a 2-layer GCN architecture, incorporating dropout and two GraphNorm layers which are applied to the output of each GCN layer, along with Leaky ReLU activation functions. Typically this GCN would be trained using a Mean Squared Error (MSE) loss, but we modify this loss function to steer our GCN towards dynamic predictions made by the dynamics-based model.
The adjusted loss function is given by:
@@ III-D3 Physics Informed GCN (PI-GCN)
where N is the number of countries, $\hat{y}_{i}$ is our model’s output, $y_{i}$ is the ground truth, and $p$ is a hyperparameter ranging from 0 to 1, signifying the influence of $\dot{y}_{i}$, the output from the dynamics-based model. This modification ensures that, during training, the model is penalized more when its prediction deviates from the linear model’s prediction, providing a form of dynamics-informed regularization. The best $p$ value was found to be 0.1 by grid search.
@@ III-E Training Procedure
3:     $model\_initial\_weights$leftarrow random
@@
9:        $model$leftarrow initialize with $model\_initial\_weights$
10:        $epochs$leftarrow 100
11:        $early\_St$leftarrow Early Stopper(patience=3)
@@
13:        $best\_model\_weights$ $\leftarrow early\_St$
@@
All training and validation of the models were performed retrospectively, as outlined in Algorithm 2. This means that the model is trained and validated at each time step using only the ”observed” data available up to that time step, denoted as $d$. For each variant, the retrospective algorithm iterates through all the biweekly data associated with that variant, training the model only on the weeks that have already passed (i.e., the observed data). Note that the pre-processing is also done retrospectively, accounting for any smoothing, calculation of the $S$ values, or interpolation in the case of the dynamics-based model. This ensures that the approach is applicable in a prospective setting, where no data from the future is available.
The train/validation split is also done temporally to ensure that validation is performed only on the most recently observed data. This aims to achieve a better fit to the latest data, enhancing the model’s predictive performance. An Early Stopper monitors the validation loss and halts the training when overfitting is detected, signified by a constant and substantial increase in the validation loss. The Early Stopper saves the best-performing weights, which are reused in the next iteration. This practice eliminates the need for the model to start training from scratch, leveraging knowledge gained from previous iterations.
We use MSE, or its adjusted version, as the loss function. Another crucial aspect is how to handle data related to variants that either do not appear in a given country or appear but fail to become dominant, i.e. rapidly diminish. In such cases, where regression targets would be undefined (potentially infinite), we address this challenge by creating a mask. This mask is employed to conceal nodes corresponding to these specific variants/country pairs, ensuring they do not contribute to our training loss. Essentially, our graph models are not trained on these nodes.
@@ IV-A1 Covariant Data
CoVariants [11] is our primary dataset which provides the current state of variants and mutations of interest for SARS-CoV-2. This data is enabled by GISAID [34] and ranges from August 2020 through October 2023, with a bi-weekly resolution for 36 variants starting from 20A.EU1 to 23F.Omicron, which is current as of this work. CoVariants follows the NextStrain Clade schema [35], in which variants can descend from other variants.
@@ IV-A2 Border Restrictions Data
We considered the border restrictions data provided in the OxCGRT [31] dataset, with details on what level of border controls were enacted by governments in real-time from Jan 1, 2020, to Jan 8, 2023. OxCGRT measured the variations in government responses using their COVID-19 Government Response Stringency Index, which is a simple additive score of nine indicators ranging from school closures to vaccination policies. Of these indicators, we picked international travel closure controls since they best represent the goal of this paper in capturing the emergence and cross-border transmission of variants. These values are encoded in an ordinal scale as follows: 0 - no measures, 1 - screening, 2 - quarantine on high risk region, 3 - ban on high risk regions, 4 - total border closure. The recorded values were then scaled down to a range between 0.1 and 1, with 1 indicating no border restrictions and 0.1 representing complete border closure. Note that 0.1 was chosen over 0, acknowledging the possibility of some mobility between connecting countries due to necessary trade or border crossings via open land routes. Furthermore, given that our CoVariants data extends beyond January 2023, we extend OxCGRT by making the assumption that all border restrictions have remained unchanged since Jan 8, 2023.
@@ IV-A3 Country Graphs
We considered two graphs representing the interconnection between countries. This dataset depicts country neighborhoods according to the Correlates of War (CoW) database. Countries were also deemed neighbors if they share a land or river border or are located within 24 miles of each other across bodies of water. Adjacency: We used the country adjacency dataset [36], which offers a list of countries and their neighboring countries. OpenFlights: This is a collection of regular flights along with their routes mapped from airports around the world [37]. Given the scope of the current work, we primarily focus on airports and the flight route data, and was mainly used during ablation studies. Airports data is a UTF-8 encoded collection of over 14,000 airports, train stations and ferry terminals across the world as of January 2017. It consists of information like city/country, IATA codes, geographic coordinates and the type of the airport (air terminal, ferry station, etc.). Routes data provides information about 67663 directional routes between 3321 airports on 548 airlines. This dataset was used in an ablation study to validate the effectiveness of utilizing these flight routes as edges in our graph creation process.
A specific adjustment in preprocessing was necessary for the OpenFlights dataset, involving the mapping of city routes to their respective country routes. Leveraging the provided airport data, which contains comprehensive city and country details, facilitated this mapping process. The transformation ensured that subsequent analyses operated consistently at the country level.
@@ IV-B Benchmark Pipeline
Since we are the first to attempt solving this problem, we provide a benchmarking tool for the community. Figure 5 presents the comprehensive pipeline employed for retrospective training, validation, and testing of diverse models. Its design prioritizes user-friendliness, allowing seamless integration of new PyTorch models or direct utilization of scikit-learn models through a configuration file. Users only need to specify whether the model requires graph data, leaving the pipeline to handle the rest. Additionally, the data pre-processing is encapsulated in a separate function, providing users with the flexibility to make modifications.
@@
In the context of delay prediction, considering the biweekly data, the target is consistently a multiple of 14 days. Therefore, the output of our models is always rounded up to the nearest multiple of 14. Furthermore, evaluations exclude countries where a variant has already become dominant or never achieves dominance.
IV-C Pre-processing
The initial pre-processing stage aligned the country names across diverse datasets, with an emphasis on inclusion based on data availability. Only countries present in all datasets were retained, ensuring a standardized set for subsequent analyses. For each country, the prevalence of each variant was calculated at each timestep. Variants with a prevalence of less than 5% were disregarded in our analysis. To calculate the prevalence ratios, the two variants with the highest prevalence were identified for each timestep. For each variant on that particular day, the prevalence ratio was calculated. Finally, variants that did not reach a prevalence ratio of 0.2 or those that were present in less than three countries were filtered out. These steps ensured that only variants with a significant presence and wide geographic distribution were included in the subsequent analysis. All subsequent pre-processing steps were conducted retrospectively. This implies that as time progresses, more data is used for computing the global growth rates ($S$) and for training.
IV-D Benchmark Evaluation Metrics
To measure performance, we define Mean/Median Absolute Errors (MAE/MedAE) as follows:
@@ IV-D Benchmark Evaluation Metrics
IV-E Benchmark Models